The addition of hydrochloric acid to a silver nitrate solution precipitates silver chloride according to the following reaction:


AgNO3(aq) + HCl(aq) → AgCl(s) + HNO3(aq)


When 50.0 mL of 0.100 mol L−1 AgNO3 is combined with 50.0 mL of 0.100 mol L−1 HCl in a coffee-cup calorimeter, the temperature changes from 23.40 °C to 24.21 °C. Calculate ΔrH for the reaction as written. Use 1.00 g mL−1 as the density of the solution and C = 4.184 J g−1 °C−1 as the specific heat capacity.


1
Expert's answer
2021-02-09T03:47:28-0500

Solution:


q = mc (T2 - T1)

q = 100g (4.184 "(\frac{J}{gC})") (24.21 - 23.40)C

q = 301 J

Moles AgCl formed = 0.05 L * 0.100 "(\frac{mole}{J})" = "5*10^{-3}mole"

"\Delta H_{reaction}" = "\dfrac{-301J}{5*10^{-3}mole}=6*10^{4}(\frac{J}{mole})=6\;(\frac{kJ}{mole})"